OpenStudy (anonymous):

Substitution Method 3x – 2y = 8 x = 2 please show me steps and check answer steps

OpenStudy (smarty):

Substitute the second part x=2 into the first one so. 3(2)-2y=8 6-2y=8 -2y=2 y=-1

OpenStudy (smarty):

Then plug -1 in for y and solve for x so x=2 as 3x-2(-1)=8 3x+2=8 3x=6 x=2

OpenStudy (anonymous):

your awesome wish it was that easy for me

OpenStudy (smarty):

No worries :D

OpenStudy (anonymous):

thank u i apparate everyone that has helped me with this course